OpenStudy (anonymous):

evaluate h^3 when h=6 is the answer h^3=18

OpenStudy (dumbcow):

6^3 = 6*6*6 = 216

OpenStudy (anonymous):

no remember you take the number times the amount. 6^3=6*6*6=?

OpenStudy (anonymous):

thnks dumcow I had that answer at first and then changee it